OpenStudy (anonymous):

what is the answer to x2+10x+5=0? please help me im so confused!

OpenStudy (anonymous):

x=0

OpenStudy (zrsmith):

Using the quadratic equation x= -5-2*sqrt(5) = 2*sqrt(5)-5 x=-5-2*sqrt(5)

OpenStudy (zrsmith):